### v1

@@ 2.3 Opinion Consensus with Messengers
In order to prevent the system from fragmenting into isolated opinion clusters, i.e. echo chambers, we extend the model by introducing a Messenger state. In this state, each agent behaves as data moving randomly in space. Messengers can make the information propagate and spread beyond the limitation of the communication range. A key outcome of introducing Messengers is the achievement of opinion consensus even in low connectivity regimes, which was previously impossible. However, this also introduces a new dimension to the problem: deciding which behavior agents should follow. As we explain in Sec. 4.1, agents switch between the two behaviors following a DMP. Here, we study how the two main parameters of the DMP ($p_{\text{M}}$ and $p_{\text{E}}$), the transition probabilities, modulate the collective performance in terms of opinion consensus.
We illustrate the normalized opinion precision error ($E_{\text{P}}^{\text{O}}$) across the parameter space in Fig. 5. Each point in the 2D space of $p_{\text{M}}$ and $p_{\text{E}}$ represents a pair of parameters corresponding to a specific configuration for the DMP. The bottom-right corner point is analogous to the baseline setup, where the ratio of Messengers is zero. This serves as our reference for normalization, with each point’s normalized performance being the ratio of its absolute performance to that of the baseline. The figure shows that the baseline setup (bottom-right corner), where all agents are Exploiters, is not optimal. The higher performance of other regions indicates that introducing the Messengers can promote the consensus of opinions. Also, the collective performance is a significant function of the DMP parameters. Changing the parameters can move the system across different regimes, with certain configurations being locally optimal. Identifying these local optimal regions will aid in understanding the conditions necessary for achieving consensus, and provide guidelines for designing such systems. For ease of reference, we marked the parameter space with specific regions of qualitatively similar behavior.
@@
According to the properties of the DMP (see Sec. 4.3), the bottom and left parts of the parameter space correspond to the regions with slow switching (large $\tau_{\text{S}}$) of states on average, with agents remaining in at least one of the states for very long times. Given that agents in these regions rarely switch their states (see Fig. 4-b), we can consider them as specialized individuals. This is in contrast to the top-right corner, where generalized agents frequently switch between the two states, corresponding to short sojourn times. The generalized agents experience both states during the course of the experiment. Additionally, if we divide the space off-diagonally into two triangles, the lower right triangle predominates with Exploiters. We already know that the ratio of Messenger population decreases from the top-left to the bottom-right corner of the parameter space (see Fig. 4-c.)
@@ 2.3.1 Integration-vs-Information Tradeoff
In Fig. 5, as we go up diagonally from the bottom-right corner to the upper-left corner, we encounter collectives with varying proportions of Exploiters. For instance, in region R1, Exploiters constitute the largest share of the population on average resulting in overly clustered collectives, where agents immediately entrap to the local echo chambers. The opinion consensus performance in this region is deficient due to the prevalence of many greedy Exploiters. We consider an Exploiter as an integrator of information by processing it following the social learning DeGroot model (see Sec. 4.1.) Therefore, R1 is dominated by agents acting as an integrator of the available information. Their greedy and exploitative behavior prevents the flow of new information across clusters. In this setting, the system exhibits the most conservative behavior in terms of mobility and tends to over-exploit in terms of information aggregation. For engineering applications, especially where moving is costly or hazardous, selecting parameters closer to this region might be preferred. Conversely, R3 represents an overly spread region where extreme exploration of Messengers and inadequate processing of available information negatively impact performance. In this region, Messengers, acting as moving information, are redundant; while there are insufficient Exploiters to integrate their information into the collective. In R3, collectives demonstrate the highest network plasticity, meaning any two random agents are likely to meet each other, irrespective of their opinions. R1 and R3 are extremes of either exploitation or exploration, respectively. The high-performing regions are situated between these two extremes, closer to the off-diagonal line (the blue dashed line). We labeled this balanced region as R2. This region contains all local optimal regions between the blue and red contours. Next, we examine different locally optimal regions within R2, focusing on varying switching timescales.
@@ 2.3.2 Generalization-vs-Specialization Tradeoff
We know that moving from the bottom-left to the top-right corner of the parameter space decreases the timescale (increases the frequency) of switching between the two states, as shown in Fig. 4-b. This change in the DMP timescale influences the optimal regions within R2. Near the bottom-left corner, R2-a is located below the off-diagonal line. The corner represents specialized agents who do not switch their state during the simulation and maintain their initial Exploiter or Messenger states. The tail of R2-a aligns with the iso-lines of Messenger ratios (see Fig. 4-c). This indicates that the system requires a specific ratio of Messengers that is significantly less than half. Accelerating the switching frequency (towards the top-right) increases the likelihood of agents experiencing both states, thus making them more generalized. As we go up in this space, the high-performing points gradually shift towards higher Messenger ratios near the center of the parameter space. This pattern suggests that, on average, more Messengers are needed in generalized collectives compared to the specialized ones. Increasing the switching frequency further results in the branching of R2 into two distinct sub-regions: R2-b and R2-c.
R2-b, the lower branch of this fast-switching region, is not parallel to the iso-lines of $m$ (Messenger ratio). This implies that the optimality of R2-b is caused by both properties of the DMP. In this region, the points below the off-diagonal line (lower $m$) are associated with longer timescales. Whereas points with higher Messenger ratios (above the line) favor faster switching. Longer timescales for Messengers mean that they have more time to transport the information they carry, effectively lengthening the link they create. However, compared to R2-a, Messengers in R2-b do not traverse for long durations. Consequently, R2-b is an optimal region, especially in scenarios where movement and exploration incur significant costs. Observations of agents’ behavior during simulations (see the Supplementary material) reveal that this configuration enhances the collective’s exploration behavior. Unlike R2-b, R2-c comprises a majority of Messengers and is elongated parallel to the iso-lines of Messenger ratios. This pattern suggests that the optimal parameter configurations in R2-c primary depend on $m$, and are not strongly correlated with $\tau_{\text{s}}$. This region is distinguished by effective information mixing, promoted by the abundance of Messengers moving randomly in space. The slower configurations outside of R2-c are less optimal, as Messengers transport information across longer distances than necessary. The superiority of faster dynamics in this region underscores the significance of spatio-temporal coupling in the problem. We later show that, compared to R2-b, convergence is achieved more rapidly in the R2-c region (see Fig. 6.)
@@ 2.3.3 Temporal Evolution of Parameter Space
By examining snapshots of the parameter space at different time steps, we obtain insights into the temporal evolution of the various regimes in the parameter space. Observing the system’s progression over time helps us to track the emergence and development of each of the optimal regions (as shown in Fig. 6). This approach enables a comparison of the temporal characteristics of each region and allows us to evaluate their performance in scenarios with different time budgets available for the overall task. For instance, the high-performing optimal region located at the bottom-left of the space demonstrates consistent performance, irrespective of the time limit. It maintains its optimality across all time limits, suggesting a more robust and reliable performance. This region is characterized by agents that are specialized in their initial states. The gradual slight diagonal shift of this region’s tail to the lower right, observable from Fig. 6-a to e, indicates that a lower (specialized) Messenger ratio is favorable in scenarios with longer time budgets.
In contrast, the generalized Exploiters, located in the top right narrow valley (R2-b), emerge only when the system has enough time to process the task. It is important to note that this region is advantageous as it minimizes precision error in both the physical and opinion domains, as we discussed in Sec. 2.4. Additionally, we observe a bifurcation in the fast-switching parameter regime. Specifically, the small head of the elongated R2 (see Fig. 6-b) grows over time and eventually branches into two distinct regions (R2-b and R2-c). The gap between the two branches widens over time. Another notable temporal shift occurs in the generalized Messenger region, the larger upper left branch (R2-c) by gradually ascending toward higher Messenger ratios. This suggests that in scenarios with higher time budgets, increasing exploration efforts–by assigning more agents to perform random walks–is advantageous. The final aspect to highlight is the continual improvement in the precision of the optimal regions. The gained precision over time underlines a speed-accuracy tradeoff in this scenario. In contrast, the extreme parameter setups, such as those characterized by excessive exploitation or randomness (R1 and R3), do not exhibit any significant improvement over time.
@@ 2.4 Contour Capturing Performance
We turn now to evaluating the precision error in the spatial domain, measured by $E_{\text{P}}^{\text{S}}$. This metric quantifies the collective performance in contour-capturing tasks and is particularly useful when access to the internal opinion of agents is not feasible. The results, as depicted in Fig. 7-a or b, do not offer any additional high-performing regions compared to those already identified in Fig. 5. However, the local optimal regions above the diagonal disappear. This is because excessive random walks by too many Messengers do not contribute to spatial convergence. The difference between the two metrics, $E_{\text{P}}^{\text{S}}$ and $E_{\text{P}}^{\text{O}}$, signifies that while consensus may be achieved in the opinion domain, it does not necessarily translate to a spatial consensus.
@@ 2.4.1 The Effect of Initial State Distribution of the DMP
We already have established that the temporal properties of the DMP significantly affect collective performance. Another important aspect to consider is the distinction between the system’s transient and stationary behaviors. Up to this point, we have assumed that the DMP begins with the expected analytical ratio of Messengers, thereby initially placing the process in its stationary state. However, any deviation from this stationary state puts the system into a transient phase. The progression towards reaching the stationary state is time-dependent and the transient behavior of the DMP also influences the collective performance. Specifically, we show that the transient behavior is linked to the DMP relaxation time ($\tau_{\text{C}}$)  [42] (see Fig. 10 in the appendix). Next, we will explore how the system’s behavior is affected by its initial condition, particularly focusing on the initial population of Messengers. To show this difference in Fig. 7, we conduct a comparison to a case where the system starts without any Messengers. We know that, given fixed transition rates, the system requires time to attain the expected analytical stationary properties, namely the Messenger population ratio.
The observed differences are particularly noticeable in DMPs with large timescales (bottom-left corner in Fig. 7). A comparison of the two figures shows that for slow-switching dynamics, the tail of the optimal region bends upward under non-stationary initial conditions. This suggests that the absence of initial Messengers is compensated for with higher $p_{\text{M}}$ values. In such scenarios, the actual average Messenger ratio falls below its expected stationary value due to the transient behavior of the Markov process.
Same as the temporal development studied earlier in Fig. 6, by comparing the parameter space at different time steps we observe that systems with different time-scales reach their stationary state with different speeds. For example, systems with fast-switching dynamics in the top-right corner reach their stationary states more quickly, which makes them more robust to varying initial conditions within the same time constraints. In contrast, slower systems require more time to gradually approach their stationary states. Consequently, we observe that the horizontal tail of R2 shifts downward, eventually aligning with a shape similar to that depicted in Fig. 6-a. With these observations, we highlight the importance of accounting for transition effects, particularly in scenarios where the properties of the system must dynamically change from one configuration to another. They also provide insights for engineers in designing robust systems against variations of initial conditions and selecting parameters from different optimal regions.
@@ 2.4.2 Different Information Distributions
Figure 8: The performance of the collective in different environment distributions. First row) the information distribution in space, and its mean contour line (red dashed line). The cyan dots show the scattered position of agents at the end of each simulation for the specific parameter marked by the green dot in the respective figure in the second row. Second row) the precision error ($E_{\text{P}}^{\text{S}}$) for the corresponding environment.
Our main focus in this paper is on the specific case where the spatial distribution of the information is radial, i.e. the iso-contour lines form circular shapes. This configuration resembles a range of natural situations where a point-like source isometrically emits a concentration into its surroundings, which may then spread through processes like diffusion or convection. To ensure generality, we also tested our model’s performance in environments with various other distributions. The first row of Fig. 8 presents different information distributions, where agents (depicted as cyan dots) converge to the zero-bias (ground-truth) contour line (indicated by a red dashed line). The introduction of Messenger with tuned DMP parameters shows an improved performance across all four environmental benchmarks.
In the second row of Fig. 8, we present the optimal regions within the parameter space for different information distributions. While the shape of these optimal regions varies slightly depending on the distribution, the principles and the arguments discussed in other subsections of the paper remain applicable to each of them. For this subsection, we restricted our analysis to the top-right quarter of the previously broader parameter space, since the most non-trivial optimal regions are located in this quarter. Furthermore, we selected a specific optimal parameter pair (marked with the green dot) to demonstrate the final agent spatial distribution for this particular DMP configuration. This visualization confirms that the chosen parameter set results in a satisfactory spatial consensus across all tested information distributions. Additionally, we adapt our model to an abstract 1-dimensional environment, with a few modifications, such as the implementation of random walk (see appendix). The outcomes from this setting show only negligible qualitative differences compared to the 2D environment.
@@ 3 Discussion
The integration of explicit spatial behavior into the collective opinion dynamics model introduces a new dimension to the complexity of the model. We demonstrated this through several components: a) initializing opinions based on spatial distribution, b) defining the interaction network according to spatial local proximity, and c) driving agents’ movement in space through homophily. The interplay of these elements results in complex collective behavior, where homophily in space leads to emergent patterns that match with the spatial distribution of information, the so-called ‘contour capturing behavior.’ We interpret this collective pattern formation as a type of spatial consensus and show that the consensus on the pattern is conditional on having sufficient network connectivity. Our simulations for collectives with low connectivity indicate that the greedy behavior of agents leads to the collective becoming trapped in local optima, a process we associate with the emergence of echo chambers. A unique feature of these echo chambers is their similar shapes, which mirror the spatial information distribution. The spatial patterns constructed by the echo chambers reflect the properties of the information landscape and are a consequence of the spatial dynamics of the system, typically ignored in opinion formation models. Also, our findings indicate that these echo chambers inhibit the flow of information, which prevents the collective’s ability to reach a precise consensus. It is important to clarify that the impact of network connectivity is relative to the size of agents’ distribution. Low connectivity can stem from either a limited communication range or an extensive initial diversity in opinion space. This observation highlights a potential disadvantage of excessive opinion diversity in a collective when the communication range is limited.
@@
We suggested that a method to overcome the local traps inherent in low-connectivity networks is to indirectly extend the effective communication range. This is particularly relevant in systems comprising mobile agents. By leveraging mobility and employing certain individuals as embodied data carriers, individuals can interact beyond the physical limits of their communication. To this end, we introduce a new behavioral role for agents, the Messenger state. This state allows designated agents to function exclusively as carriers of information. A Messenger is a stubborn agent with a fixed opinion, sharing its opinion while moving randomly in space. These agents can be viewed as links to the history of opinions of agents, with their independent random movement promoting broader exploration. Yet, defining a new state for agents poses a new challenge: determining how agents should select their state. To address this challenge, we employed the Dichotomous Markov Process (DMP) as a distributed switching mechanism for agents to switch between behaviors. DMP gives agents a mechanism to decide at each time step whether they should switch to the other state or stay in their current state. A key advantage of this approach is its decentralization and scalability.
@@
Still, the two parameters of the DMP add a dimension of freedom to the problem that requires consideration. Modifying the DMP parameters directly determines the temporal properties of the switching mechanism for a single agent. This, in turn, indirectly modifies the collective’s properties, such as the ratio of Messengers and the speed of switching (denoted as $m$, and $\tau_{\text{S}}$). The expected value of these two properties can be described analytically as a function of the DMP parameters. We used these analytical values to validate the implementation of the Markov process in our numerical simulations. We showed that these properties significantly influence the collective performance, in achieving consensus in both the opinion and space domains.
@@
Our numerical results mark several local optimal regions where the collective achieves the highest precision in consensus, which is the ultimate objective in this scenario. Each region exhibits unique characteristics, reflecting a spectrum of collective behaviors derived from the DMP. We distinguished these regions based on the two key properties of the DMP. The high-performing regions for achieving consensus vary depending on the consensus domain–opinion or spatial–used to evaluate the system. By adjusting the DMP parameters, agents can adapt their performance and navigate the various tradeoffs we have identified. We elaborate on some of these tradeoffs such as information versus integration, and generalization versus specialization. Generally, extremes in the number of Messengers–either too few or too many–prove suboptimal. Pushing the system towards them results in either excessive exploitation and greedy behavior or randomness and chaos. The high-performing settings were located in the middle ground. We have categorized these regions into three distinct groups, each distinguished by the switching frequency of the DMP.
@@
The range of switching frequency spans from no switching to excessively fast switching. Agents at these extremes are identified as being specialized or generalized, respectively. Compared to the Messengers ratio, the dimension of switching frequency introduces a more nuanced trade-off, as we elaborated in the Results section. We found that even a minimal number of Messengers can reestablish consensus in collectives when agents maintain fixed states throughout the simulation–being specialized. On the other hand, faster switching can result in two distinct high-performing behaviors: improved exploration for lower Messenger ratios; or conservation of information for higher Messenger ratios and updating the opinions only rarely. In the former, the fast-switching Messengers perform random walks long enough to help the collective escape the local optima of echo chambers. The latter is the case for collectives mainly composed of Messengers, who update their opinion quite rarely and integrate information only for a short duration. However, this duration is sufficient for them to process information and reach a consensus.
@@
We also highlighted the transient and temporal aspects of collective behavior caused by the DMP. This perspective gives insight into understanding and studying the robustness of the performance against non-stationary behavior or dynamic conditions, particularly in scenarios where the initial conditions of the system are different than the expected stationary properties. For example, when the system starts with no Messengers and gradually increases the Messenger population to reach its expected value. Such analyses help engineer the system and guide designers to select parameters that are more robust against transient behaviors. Moreover, we simulated the model in various information distributions with both uni- and multi-modal shapes to assess the dependency of behavior on specific cases. The results showed negligible qualitative differences across different information distributions.
@@
While only varying the DMP parameters and keeping the other parameters of the model constant has allowed for an in-depth exploration of the system, future research could expand on this by investigating the influence of the other parameters of the model. This includes aspects such as collective density, arena size, or information noise. Also, investigating the one-dimensional version, which resembles classical opinion dynamics more closely, could provide valuable insights into a range of related issues in the field.

### v2

@@ 2.3 Opinion Consensus with Messengers
The ability to arrive at the consensus for the heterogeneous collective consisting of both Messengers and Exploiters depends on two parameters of the DMP ($p_{\text{M}}$ and $p_{\text{E}}$) governing the switching process. A balanced set of properties ($m$ and $\tau_{\text{S}}$) is required to enable Messengers to propagate information efficiently beyond the limitation of the communication range. This also applies to the Exploiters, which are needed to process the information Messengers carry.
We illustrate the normalized opinion precision error ($E_{\text{P}}^{\text{O}}$) across the parameter space in Fig. 5. Each point in the 2D space of $p_{\text{M}}$ and $p_{\text{E}}$ represents a pair of parameters corresponding to a specific configuration for the DMP. The bottom-right corner point is analogous to the baseline setup, where the ratio of Messengers is zero. This serves as our reference for normalization, with each point’s normalized performance being the ratio of its absolute performance to that of the baseline. The figure shows that the baseline setup (bottom-right corner), where all agents are Exploiters, is not optimal. The higher performance of other regions indicates that introducing the Messengers promotes consensus. Changing the parameters can move the system across different regimes, with certain configurations being locally optimal. The analysis of these local optimal regions increases our understanding of the conditions necessary for achieving consensus and provides guidelines for designing such systems. For ease of reference, we labeled different regions of the parameter space exhibiting qualitatively similar behavior.
@@
According to the properties of the DMP (see Sec. 4.1.4), the bottom and left parts of the parameter space correspond to the regions with slow switching of states on average (large $\tau_{\text{S}}$), with agents remaining in at least one of the states for very long times. Given that agents in these regions rarely switch their states (see Fig. 4-b), we can consider them as specialized individuals. This is in contrast to the top-right corner, where generalized agents frequently switch between the two states, corresponding to short sojourn times. The generalized agents experience both states during the course of the experiment. Additionally, if we divide the space off-diagonally into two triangles, with the lower right triangle dominated by Exploiters.
@@ 2.3.1 Integration-vs-Information Tradeoff
In Fig. 5, as we go up diagonally from the bottom-right corner to the upper-left corner, we encounter collectives with varying proportions of Messengers. For instance, in region R1, Exploiters constitute the largest share of the population on average resulting in strongly clustered collectives, where agents immediately get trapped in local echo chambers. The opinion consensus performance in this region is deficient due to the prevalence of many greedy Exploiters. We consider an Exploiter agent to integrate and process information according to the DeGroot model for social learning (see Sec. 4.1.) Therefore, R1 is dominated by the integration of the available information. Their greedy and exploitative behavior prevents the flow of new information across clusters. In this setting, the system exhibits the most conservative behavior in terms of mobility and tends to over-exploit in terms of information aggregation. For engineering applications, especially where moving is costly or hazardous, selecting parameters closer to this region might be preferred. Conversely, R3 represents an overly dispersed region where extreme exploration of Messengers and inadequate processing of available information negatively impact performance. In this region, too many Messengers, acting as moving information, are redundant; while there are insufficient Exploiters to integrate new information into the collective. In R3, collectives demonstrate the highest network plasticity, meaning any two random agents are likely to meet each other, irrespective of their opinions. R1 and R3 are extremes of either exploitation or exploration, respectively. The high-performing regions are situated between these two extremes, closer to the off-diagonal line (the blue dashed line). We labeled this balanced region as R2. This region contains all local optima between the blue and red contours confining R1 and R3, respectively. Next, we examine different locally optimal regions within R2, focusing on varying switching timescales.
@@ 2.3.2 Generalization-vs-Specialization Tradeoff
Moving from the bottom-left to the top-right corner of the parameter space decreases the timescale (increases the frequency) of switching, as shown in Fig. 4-b (see also Eq. 7). The change in the DMP timescale influences the optimality of regions within R2. Near the bottom-left corner, R2-a is located below the off-diagonal line. The corner represents specialized agents, that effectively, due to the low switching probabilities, do not switch their state during the simulation and maintain their initial Exploiter or Messenger states. The tail of R2-a aligns with the iso-lines of Messenger ratios (see Fig. 4-c). This indicates that the system with static roles performs best for a specific ratio of Messengers that is significantly less than half. Accelerating the switching frequency (towards the top-right) increases the likelihood of agents experiencing both states, thus making them more generalist. As we go up in this space, the high-performing points gradually shift towards higher Messenger ratios near the center of the parameter space. This pattern suggests that, on average, more Messengers are needed in collectives of generalists compared to the specialized ones. Increasing the switching frequency further results in the branching of R2 into two distinct sub-regions: R2-b and R2-c.
R2-b, the lower branch of this fast-switching region, is not parallel to the iso-lines of $m$. This implies that the optimality of R2-b is caused by both properties of the DMP. In this region, the points below the off-diagonal line (lower $m$) are associated with longer timescales. Whereas points with higher Messenger ratios (above the line) favor faster switching. Longer timescales for Messengers mean that they have more time to transport the information they carry, effectively lengthening the link they create. In other words, smaller numbers of Messengers should move for longer durations. However, compared to R2-a, Messengers in R2-b switch back to the Exploiter state much faster. Consequently, R2-b is an optimal region, especially in scenarios where long random movements incur significant costs. Observations of agents’ behavior during simulations reveal that this configuration indirectly enhances the collective’s exploration behavior.
Unlike R2-b, R2-c comprises a majority of Messengers and is elongated parallel to the iso-lines of Messenger ratios. This pattern suggests that the optimal parameter configurations in R2-c primary depend on $m$, and are not significantly sensitive to $\tau_{\text{s}}$. This region is distinguished by conservative information mixing, promoted by the abundance of Messengers moving randomly in space. The slower configurations outside of R2-c are less optimal, as Messengers transport information across longer distances than necessary. The superiority of faster dynamics in this region underscores the significance of spatio-temporal coupling in the problem. We later show that, compared to R2-b, convergence is achieved more rapidly in the R2-c region (see Fig. 6.)
@@ 2.3.3 Temporal Evolution of Parameter Space
By examining snapshots of the parameter space at different time steps, we obtain insights into the temporal evolution of the various regimes in the parameter space. Observing the system’s progression over time helps us track the emergence and development of each of the optimal regions (as shown in Fig. 6). This approach enables a comparison of the temporal characteristics of each region and allows us to evaluate their performance in scenarios with different time budgets available for the overall task. For instance, the high-performing optimal region located at the bottom-left quarter of the space demonstrates consistent performance, irrespective of the time limit. It maintains its optimality across all time limits, suggesting a more robust and reliable performance. This region is characterized by agents that are specialized according to their initial states. The gradual slight diagonal shift of this region’s tail to the lower right, observable from Fig. 6-a to Fig. 6-e, indicates that a lower (specialized) Messenger ratio is favorable in scenarios with longer time budgets.
In contrast, the local optimum for generalist Exploiters, located in the top right narrow valley in the precision error (R2-b), emerges only when the system has enough time to process the task. It is important to note that this region is advantageous as it minimizes precision error in both the physical and opinion domains, as we discuss further below in Sec. 2.4. Additionally, we observe a bifurcation in the fast-switching parameter regime. Specifically, the small head of the elongated R2 (see Fig. 6-b) grows over time and eventually branches into two distinct regions (R2-b and R2-c). The gap between the two branches widens over time. Another notable temporal shift occurs in the region with generalist agents and a time-averaged majority of Messengers, the larger upper left branch (R2-c,) which gradually ascends toward higher Messenger ratios. This suggests that in scenarios with higher time budgets, increasing exploration efforts–by assigning more agents to perform random walks–is advantageous. The final aspect to highlight is the continuous improvement in the precision of the optimal regions. The precision gain over time demonstrates the existence of a speed-accuracy tradeoff in this scenario. In contrast, extreme parameter setups, such as those characterized by excessive exploitation (R1) and random diffusion of Messenger (R3), do not exhibit any significant improvement over time.
@@ 2.4 Contour Capturing Performance
We turn now to evaluating the precision error in the spatial domain, measured by $E_{\text{P}}^{\text{S}}$. This metric quantifies the collective performance in contour-capturing tasks and is particularly useful when access to the internal opinion of agents is not possible. The results, as depicted in Fig. 7-a or Fig. 7-b, do not offer any additional high-performing regions compared to those already identified in Fig. 5. However, the local optimal regions above the diagonal disappear. This is because excessive random walks by too many Messengers do not contribute to spatial convergence. The difference between the two metrics, $E_{\text{P}}^{\text{S}}$ and $E_{\text{P}}^{\text{O}}$, signifies that while consensus may be achieved in the opinion domain, it does not necessarily translate to a spatial consensus.
@@ 2.4.1 The Effect of Initial State Distribution of the DMP
We already have established that the temporal properties of the DMP significantly affect collective performance. Another important aspect to consider is the distinction between the DMP transient and stationary behaviors. Up to this point, we have assumed that the DMP begins with the expected analytical ratio of Messengers, thereby initially placing the process in its stationary state. However, any deviation from this stationary state puts the system into a transient phase. The progression towards reaching the stationary state is time-dependent and the transient behavior of the DMP also influences the collective performance. Specifically, we show that the transient behavior is linked to the DMP relaxation time ($\tau_{\text{C}}$)  [46] (see Fig. 10 and Eq. Relaxation Time of the DMP in the Appendix). Next, we will explore how the system’s behavior is affected by its initial condition, particularly focusing on the initial population of Messengers. To show this difference in Fig. 7, we conduct a comparison to a case where the system starts without any Messengers. We know that, given fixed transition rates, the system requires time to attain the expected analytical stationary properties, namely the Messenger population ratio.
The observed differences are particularly noticeable in DMPs with slow timescales (bottom-left corner in Fig. 7). A comparison of the two figures shows that for slow-switching dynamics, the tail of the optimal region bends upward under non-stationary initial conditions. This suggests that the absence of initial Messengers is compensated for with higher $p_{\text{M}}$ values. In such scenarios, the actual time-averaged Messenger ratio is below its expected stationary value due to the transient behavior of the Markov process.
Same as the temporal development studied earlier in Fig. 6, by comparing the parameter space at different time steps we observe that systems with different time-scales reach their stationary state with different speeds (see Fig. 11). For example, systems with fast-switching dynamics in the top-right corner reach their stationary states more quickly, which makes them more robust to varying initial conditions within the same time constraints. In contrast, slower systems require more time to gradually approach their stationary states. Consequently, we observe that the horizontal tail of R2 shifts downward, eventually aligning with a shape similar to that depicted in Fig. 6-a. With these observations, we highlight the importance of accounting for transient effects, particularly in scenarios where the properties of the system must dynamically change from one configuration to another. They also provide insights into the robustness of each configuration against variations of initial conditions and selecting parameters from different optimal regions.
@@ 2.4.2 Different Information Distributions
Figure 8: The performance of the collective in different environment distributions. First row) the information landscape, and its mean contour line (red dashed line). The cyan dots show the scattered position of agents at the end of each simulation for the specific parameter marked by the green dot in the respective figure in the second row. Second row) the precision error ($E_{\text{P}}^{\text{S}}$) for the corresponding environment.
Our main focus in this paper is on the specific case where the spatial distribution of the information is radial, i.e., the iso-contour lines are circles. This configuration resembles a range of natural situations where a point-like source isometrically emits a concentration into its surroundings, which may then spread through processes like diffusion or convection. To ensure generality, we also tested our model’s performance in environments with various other distributions. The first row of Fig. 8 presents different information distributions, where agents (depicted as cyan dots) converge to the zero-bias (ground-truth) contour line (indicated by a red dashed line). The introduction of Messenger with tuned DMP parameters shows an improved performance across all four environmental benchmarks.
In the second row of Fig. 8, we present the optimal regions within the parameter space for different information distributions. While the shape of these optimal regions varies slightly depending on the distribution, the general features of the error landscape and the underlying mechanisms discussed in previous subsections remain unchanged. Here, we restricted our analysis to the top-right quarter of the parameter space, where the most non-trivial optimal regions are located. Furthermore, we selected a specific optimal parameter pair (marked with the green dot) to demonstrate the final agent spatial distribution for this particular DMP configuration. This visualization confirms that the chosen parameter set results in a satisfactory spatial consensus across all tested information distributions. Additionally, we also adapted our model to an abstract 1-dimensional environment, with a few modifications, such as the implementation of the random walk (see Appendix). The corresponding results show only negligible qualitative differences compared to the 2D environment.
@@ 3 Discussion
The integration of explicit spatial behavior into the collective opinion dynamics model introduces a new dimension to the complexity of the model. We demonstrated this through several components: a) initializing opinions based on spatial distribution, b) defining the interaction network according to spatial local proximity, and c) driving agents’ movement in space through homophily. The interplay of these elements results in complex collective behavior, where homophily in space leads to emergent patterns that trace features of the spatial distribution of information, resulting in the so-called ‘contour capturing behavior.’ This collective pattern formation represents a type of spatial consensus, and we show that it is conditional on having sufficient network connectivity. Our simulations for collectives with low connectivity indicate that the greedy behavior of agents leads to the collective becoming trapped in local optima, a process we associate with the emergence of echo chambers. A unique feature of these echo chambers is their similar shapes, which mirror the spatial information distribution. The spatial patterns constructed by the echo chambers reflect the properties of the information landscape and are a consequence of the spatial dynamics of the system, typically ignored in opinion formation models. Also, our findings indicate that these echo chambers inhibit the flow of information, which prevents the collective’s ability to reach a precise consensus. It is important to clarify that the impact of network connectivity is relative to the size of agents’ distribution. Low connectivity can stem from either a limited communication range or a large initial diversity in opinion space. This observation highlights a potential disadvantage of excessive opinion diversity in a collective when the communication range is limited.
@@
We suggested that a method to overcome the local traps inherent in low-connectivity networks is to indirectly extend the effective communication range. This is particularly relevant in systems comprising mobile agents. By leveraging mobility and employing certain individuals as embodied data carriers, individuals can interact beyond the physical limits of their communication. To this end, we introduce a new behavioral role for agents, the Messenger state. This state allows designated agents to function exclusively as carriers of information. A Messenger is a stubborn agent with a fixed opinion, sharing its opinion while moving randomly in space. These agents can be viewed as mobile “quasi-memory” of agents’ opinions, with their independent random movement promoting broader exploration. We employed the Dichotomous Markov Process (DMP) as a simple, distributed switching mechanism between behaviors. The agents decide randomly at each time step whether they should switch to the other state or stay in their current state. A key advantage of this approach is its decentralization, minimal computation requirements, and scalability.
@@
Still, the two parameters of the DMP add an additional degree of freedom. Modifying the DMP parameters determines the temporal properties of the switching mechanism for a single agent. This, in turn, indirectly modifies the collective’s properties, such as the ratio of Messengers and the average speed of switching (denoted as $m$, and $\tau_{\text{S}}$). The expected value of these two properties can be derived analytically as a function of the DMP parameters. We showed that these properties significantly influence the collective performance, in achieving consensus in both the opinion and the spatial domains.
@@
Our numerical results identify several local optimal regions where the collective achieves the highest precision in consensus, which is the ultimate objective in this scenario. Each region exhibits unique characteristics, reflecting a spectrum of collective behaviors derived from the DMP. We distinguished these regions based on the two key properties of the DMP. The high-performing regions for achieving consensus vary depending on the consensus domain (opinion or spatial) used to evaluate the system. By adjusting the DMP parameters, agents can adapt their performance and navigate the various tradeoffs we have identified. We elaborated on some of these tradeoffs such as information versus integration, and generalization versus specialization. Generally, extremes in the number of Messengers (either too few or too many) prove sub-optimal. Pushing the system towards them results in either excessive exploitation or random behavior. The high-performing settings were located in the intermediate parameter regions. We have categorized these regions into three distinct groups, each distinguished by the switching frequency of the DMP.
@@
The range of switching frequency spans from no switching to excessively fast switching. Agents at these extremes are identified as being specialists or generalists, respectively. Compared to the Messengers ratio, the dimension of switching frequency introduces a more nuanced trade-off. We found that even a minimal number of Messengers can reestablish consensus in collectives of specialists, i.e., when agents maintain fixed states throughout the simulation. On the other hand, faster switching can result in two distinct high-performing behaviors: improved exploration for lower Messenger ratios; or conservation of information at higher Messenger ratios, resulting in a slow update of opinions. In the former, the fast-switching Messengers perform random walks long enough to help the collective escape the local optima, corresponding to echo chambers. The latter is the case for collectives mainly composed of Messengers, who update their opinion quite rarely and integrate information only for a short duration. However, this duration is sufficient for them to process information and reach a consensus.
@@
We also highlighted the transient and temporal aspects of collective behavior caused by the DMP. This perspective gives insight into understanding and studying the robustness of the performance against non-stationary behavior or dynamic environments, particularly in scenarios where the initial conditions of the system are different than the expected stationary properties. For example, when the system starts with no Messengers and gradually increases the Messenger population to reach its expected value. Such analyses provide insights for the design of corresponding systems and the selection of parameters that are more robust against transient behaviors. Moreover, we simulated the model in various information distributions with both uni- and multi-modal shapes to assess the dependency of behavior on specific cases. The results showed negligible qualitative differences across different information distributions, demonstrating the generality of our results.
@@
While only varying the DMP parameters and keeping the other parameters of the model constant has allowed for an in-depth exploration of the system, future research could expand on this by investigating the influence of the other parameters of the model. This includes aspects such as collective density, arena size, or information noise. Also, investigating the one-dimensional version, which resembles classical opinion dynamics models, could provide valuable insights into a range of related issues in the field. An important addition to the analysis of these systems is under dynamic information landscapes, where the collective needs to rearrange itself and adapt to the changing environment. We expect that in such settings, the temporal properties of the system would be highlighted even more.